# Rydberg atomic polarimetry of radio-frequency fields

### v4

@@ IV Numerical simulations
where $I_{0}$ is the intensity of the probe light before traversing the distance $\ell$ through the vapour. The extinction coefficient $\alpha$ is proportional to a weighted sum over the off-diagonal elements of the density matrix that involves the combination of $g$ and $i$ -states [See (S6) of Supplementary Note 2]. To simulate the variation of $\alpha(\theta,\Delta_{\text{c}})$ over a spectrogram for a given RF dressing power, we therefore calculate the steady-state atomic density matrix under the combined influence of the optical and RF fields (see appendix C). The density matrix for each case includes all individual $|Fm_{F}\rangle$ hyperfine states of the ladder levels (cf. Fig. 2 and Fig. 3). For the type-1 ladder of Fig. 2 involving a $\mathrm{D}_{5/2}\!\!\leftrightarrow\!\!\mathrm{P}_{3/2}$ Rydberg transition, this amounts to 52 atomic states, whereas the type-II ladder of Fig. 3 capped by the $35\mathrm{D}_{3/2}\!\!\leftrightarrow\!\!36\mathrm{P}_{1/2}$ transition requires 36 atomic states. As in Ref. 34 an auxiliary “dummy state” is included to incorporate effective incoherent decay from transit-time broadening, collisions, and radiative decay from the Rydberg states. In our simulations, the RF power is the only fitted quantity with all other input parameters such as atomic properties, optical beam waists and powers being fixed by measured and tabulated values.
Figure 8 showcases our ability to reliably simulate spectrograms for both type-I and type-II ladders.
@@ V Discussion
For an ideal four-level ladder system, the AT-splitting that emerges in the EIT spectrum is directly proportional to the amplitude of the RF-field dressing the topmost two Rydberg states and the constant of proportionality is given by a combination of fundamental constants and the transition dipole moment between the Rydberg states. Because of the quasi-one-electron character of Rydberg atoms, the latter can be calculated to a very high accuracy. Early work, including Ref. 12, however, has introduced the misconception that for a type-I ladder with linear co-polarized fields, one is dealing with a single transition dipole moment between $r_{1}$ and $r_{2}$, and this has been reiterated pervasively in the literature[36, 37, 38, 39, 40, 14, 41, 42, 43, 44, 45, 46]. In fact, two transition dipole moments are in play differing by a factor of $\sqrt{2/3}\approx 0.8$ [cf. Fig. 4(a)] and, as noted in our above analysis, one will not be measuring a pure AT-doublet but a four-peaked spectrum. We illustrate this effect in Fig. 9(a) which presents the development in AT splitting when moving our RF emitter up close ($\sim 2$ cm) to the vapor cell (see appendix B). Focusing our attention to the blue-detuned branch ($\Delta_{c}>0$), we se how the split-out peak becomes lopsided [see Fig. 9(b)] for high RF-powers. If the goal is to establish an accurate electrometry standard based on a type-I ladder system, the four-peak nature should be accounted for because the splitting of each doublet-pair tunes at different rates: the inferred field amplitude will be skewed if an unresolved four-peak spectrum is interpreted as a pure doublet. Figure 9(c) shows a simulated spectrogram for a type-I system with Doppler broadening turned off from which Fig. 9(d) picks out the spectrum for parallel fields that underpins the skewing of the high power trace in Fig. 9(b).
The spectrogram Fig. 9(c) also elucidates that for linearly-polarized fields, the eigen-energy corresponding to a split-out peak does not change with $\theta$: the spectral features are straight vertical lines. While the position of a peak remains fixed, its prominence in the spectrum will, however, generally change. For two overlapping broadened peaks, this may give the appearance of single feature that moves continuously as $\theta$ changes. This effect can be observed as wobbling of the side-lobes in the type-1 column of Fig. 8.
@@
Figure 9: Electrometry using a type-I system with co-polarized fields. (a) Experimentally observed AT splitting of the $34D_{5/2}$ level centred on $\Delta_{\text{c}}=0$, The feature at $\Delta_{\text{c}}\sim-306$ MHz is the $34D_{3/2}$ level [cf. Fig. 6(c)]. Interaction with this state influences the $\Delta_{\text{c}}<0$ AT peak. (b) Development in the $\Delta_{\text{c}}>0$ AT peak with RF power increasing (top to bottom) from 10 to 80 (arb. units) in uniform steps. c Simulated spectrogram of the probe field transmission with Doppler broadening excluded. (d) Simulated probe field transmission for co-polarized fields when of Doppler broadening is absent.
@@
To be clear, it is incorrect to talk about “the” dipole moment between $D_{5/2}$ and $P_{3/2}$ levels for a linearly-polarized coupling field [47, 41, 42, 46, 48, 49, 50] because it depends on whether $m_{J}=|1/2|$ or $m_{J}=|3/2|$. It is also incorrect to resort to using the $m_{J}=|1/2|$ value [44, 37, 38], based on a rationale of all fields being $\pi$-polarized. This would be applicable in the absence of hyperfine structure of the ground state [see Fig. 10(a) and (b)], in which case the system is described by two parallel ideal four-state systems. For an atom in a specific, resolved hyperfine state we are, however, not dealing with parallel four-state systems, even when driving with co-polarized fields. This is evident from the diagrams in Figs. 2(b) and 3: moving up from any $m_{F}$-state of the ground level one counts more than three transition-arrows. Moreover, a vertical cascade corresponding to a given $m_{F}$ [Fig. 10(c) shows the example of $m_{F}=1$ from Fig. 2(b)] cannot be broken down into individual “unique contributions” [cf. Fig. 10(d)] as was recently stated in Ref. 51. The claim of Ref. 51 that the EIT scheme acts to probe the states of $r_{1}$ individually is invalidated by the fact that these states are degenerate and therefore indirectly coupled with each other by the RF-field via the states of $r_{2}$. Instead, when treating the the example of the $m_{F}=1$ cascade in the hyperfine basis, the splittings of the system probed by EIT is found by diagonalizing the entire $7\times 7$ matrix that describes the resonant RF interactions between the hyperfine states of $r_{1}$ and $r_{2}$ . The result of this procedure is illustrated in Fig. 10(e), replicating the $m_{F}=1$ dressed energies encountered in our procedure Fig. 4c where the eigenenergies are first found in $J$-basis basis after which the system is recast in the $F$-basis.
@@
The insight that the co-polarized cases of Figs. 2(b) and 3 are not represented by a collection of four-level systems means that there is a priori no reason why they should give rise to simple AT doublets. Indeed, as we have shown, the type-4 system features a prominent central peak in its spectrum. Interestingly, the two sidelobes of the type-4 triplet tunes as an ideal AT doublet with RF power with a tuning rate described by a single transition matrix element. The perfect absence of a central peak of the $\theta=0^{\circ}$ type-I system in Fig. 2(b) can be interpreted as a destructive interference between the multiple parallel pathways represented by (3). Regardless, the spectrum does not represent a simple AT doublet.
@@ VI Conclusion and outlook
In the current study, we have focused on the case of linearly-polarized optical and RF fields and shown that either of the two complementary spectroscopic signatures for type-1 and -4 ladders could serve to determine the angle of polarization for the RF field. The simultaneous high fidelity of our density matrix simulations for these complementary test cases instils confidence in our simulation framework’s ability to synthesize the probe light transmission for more complex combinations of polarization states. In the future we plan to pursue the determination of arbitrary polarization states of the incoming RF field through machine learning approaches where the polarization states of probe and coupling light are varied to acquire a sequence of spectrograms. For this purpose the data diversity [53] provided by the profoundly different response from type-1 and type-4 may provide a significant advantage. This would open up a new paradigm for quantum-metrological [54] polarimeters that at its heart is enabled by the quantisation of atomic angular momentum.
Figure 10: Level diagrams for a $J=\frac{1}{2}\frac{3}{2}\frac{5}{2}\frac{3}{2}$ system (a) and a $J=\frac{1}{2}\frac{3}{2}\frac{3}{2}\frac{1}{2}$ system (b) driven by $\pi$-transitions. Both realise two parallel ideal four level ladders (green background). (c) $m_{F}=1$ ladder of Fig. 2(b) in isolation. The system cannot be broken down into parts as shown in (d). (e) Level splittings obtained from the diagonalizing the matrix describing the RF-coupling (orange arrow) amongst the $F$-states of the Rydberg levels $r_{1}$ and $r_{2}$ when $m_{F}=1$. The different $F$-parts of $r_{1}$ and $r_{2}$ have been colour-coded to visualise their contributions to each hybridized dressed state. The diagonalization for the $m_{F}$=1 ladder yields five eigen-energies with separations and multiplicities as annotated.
@@ Acknowledgements.
M. Cloutman performed experiments with support from M. Chilcott, A. B. D., and S. O.; M. Cloutman analysed the data with support from M. Chilcott; M. Chilcott and A. E. carried out density matrix calculations; N. K., M. Cloutman, and M. Chilcott wrote the manuscript with input from all authors; N. K. conceived and supervised the project
@@ Appendix B RF dressing field
The $\sim 55$ GHz RF field dressing our Rydberg atoms was produced by a photomixing antenna (Toptica TeraBeam 1550). Two $\sim 1550\text{\,}\mathrm{n}\mathrm{m}$ fiber-coupled DFB lasers are combined to produce a beatnote with frequency $f_{\text{RF}}$ on an InGaAs photoconductive chip mounted on a bowtie-shaped antenna. The photomixing process produces linearly-polarized radiation at $f_{\text{RF}}$, which is collimated to a beam diverging at $12\text{\,}\mathrm{\SIUnitSymbolDegree}$ by a silicon lens. The frequencies of the DFB lasers are temperature tuned, allowing us to target the the $34\mathrm{D}_{5/2}\!\!\leftrightarrow\!\!35\mathrm{P}_{3/2}$ and $35\mathrm{D}_{3/2}\!\!\leftrightarrow\!\!36\mathrm{P}_{1/2}$ transitions at $56.714\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$ and $53.840\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$ respectively. The emitter is positioned $15\text{\,}\mathrm{c}\mathrm{m}$ above the vapour cell which is mounted on a wooden post to minimise metallic interference.
For the measurements presented in Figure 9(b) and (c), the emitter was positioned very close ($2\text{\,}\mathrm{cm}$) to the cell. A $0.75\text{\,}\mathrm{mm}$ pitch wire grid polarizer was mounted in between the emitter and the cell to suppress effects a complex RF field polarization [22] and ensure a linear polarization at the position of the atoms.
@@ Appendix C Density matrix calculations
Here the polarizations and Rabi frequencies of the probe, coupling and RF fields are encoded in the coupling operators $\Omega_{k}$ between manifolds, represented in (7) using an ordered basis of the $g$, $i$, $r_{1}$ and $r_{2}$ levels. The detuning blocks $\delta_{k}$ are scalar matrices because the individual states acted on by each block are not resolved. The problem is described in the $|Fm_{F}\rangle$ basis, to deal with the fact that the probe laser field resolves transitions between hyperfine states. The second, dissipative term of (6) is composed of collapse operators $L_{i}$ with corresponding decay rates $\gamma_{i}$. A representation of $L_{i}$ for radiative decay in the $|Fm_{F}\rangle$ basis is given in Supplementary Note 2. This term facilitates the inclusion of two classes of state decay: spontaneous emission, which has an operator for each possible polarization of the radiated photon, and incoherent decay, which is modelled with an axillary “dummy state” [34]. Incoherent decay includes transit-time broadening and collisional broadening of all states, but it also describes the radiative decay of the Rydberg states, which usually involves the emission of multiple photons on a decay path outside the state space of the simulation. The “dummy state” then decays rapidly, with equal probability, to each sublevel of the ground-state manifold.
@@
While the counter-propagating arrangement of coupling and probe beams compensates for a majority of the Doppler effect, we include the residual Doppler broadening of the room-temperature vapor by performing simulations with detunings that correspond to a range of atomic velocities along the probe/coupling beams, and averaging them according to the Maxwell-Boltzmann distribution. Our lock-in detection scheme is accounted for by removing a constant offset in the simulation: the resonant extinction of the probe beam without the coupling field.
@@
To arrive at (3) we must evaluate $\left|\sum_{F=1}^{4}C^{Fm_{F}}_{Jm_{J}Im_{I}}\left<F^{\prime}m_{F}^{\prime}\right|r_{0}\left|Fm_{F}\right\rangle\right|^{2}$.
@@
In this supplement, we provide additional details on computing the density matrix, and connecting it to experimental observations. In particular, we include the matrix form of the operators necessary to formulate the master equation (6), expanded over the hyperfine $|Fm_{F}\rangle$ basis.
Probe extinction
As discussed in the main text, the extinction of the probe beam through an atomic vapour follows the Beer-Lambert law, (5). The extinction coefficient $\alpha$ for a vapour of density $n$ at a transition of frequency $\omega_{p}$ is composed of off-diagonal elements of the atomic density matrix $\rho$ [A. Elliott et al., in preparation],
$\alpha=\frac{2n\omega_{p}}{c\epsilon_{0}\hbar}\sum_{g_{l}\in g,i_{m}\in i}|\vec{\mu}_{i_{m}g_{l}}|^{2}\text{Im}\left(\frac{\rho_{g_{l}i_{m}}}{\Omega_{i_{m}g_{l}}}\right),$ (S6)
with the sum running over the states of the $g$- and $i$-level manifolds. The inter-state couplings $\Omega_{i_{m}g_{l}}$, and dipole moments $\vec{\mu}_{i_{m}g_{l}}$ of (S6) are expressed below.
As described in appendix C, the density matrix is governed by the master equation of (6), with the atomic Hamiltonian (7).
@@
The coupling operators $\Omega_{k}$ for each transition take the usual form of the Rabi frequency,
@@
for the electric field $\vec{E}_{k}$, and transition dipole moment $\vec{\mu}_{k}$.
For convenience, we factor the coupling operators[63] into a scalar radial component $\Omega^{(r)}_{k}$ (the radial Rabi frequency) and an angular component, expanded over the spherical basis[64],
$\Omega_{k}=\Omega_{k}^{(r)}\sum_{q=-1,0,1}A^{(q)}_{k}u_{k}^{(q)}.$ (S8)
Physically, q=1,0,+1q=-1,0,+1 represents the change in angular momentum associated with $\sigma^{-},\pi,$ and $\sigma^{+}$ excitations, respectively. When expressed in terms of the Cartesian components of a unit vector $\hat{E}$ along the electric field direction (i.e., $\vec{E}=|\vec{E}|\hat{E}$), the coefficients $A^{(q)}$ are given by
$A^{(1)}$ $=-\frac{1}{\sqrt{2}}(\hat{E}_{x}+\text{i}\hat{E}_{y}),$ (S9)
$A^{(0)}$ $=\hat{E}_{z},$ (S10)
$A^{(-1)}$ $=\frac{1}{\sqrt{2}}(\hat{E}_{x}-\text{i}\hat{E}_{y}).$ (S11)
The elements of each matrix $u_{k}^{(q)}$ are found by repeated applications of the Wigner-Eckart theorem, with $\Omega_{k}^{(r)}$ absorbing the reduced matrix element,
@@
$\begin{split}&\left\langle SLJFmF\right|L_{k}^{(q)}|S^{\prime}L^{\prime}J^{\prime}F^{\prime}mF^{\prime}\rangle=\\ &\qquad(2L+1)(2J+1)(2J^{\prime}+1)(2F+1)(2F^{\prime}+1)\\ &\qquad\times\begin{Bmatrix}L^{\prime}&J^{\prime}&S\\ J&L&1\end{Bmatrix}^{2}\times\begin{Bmatrix}J^{\prime}&F^{\prime}&I\\ F&J&1\end{Bmatrix}^{2}\\ &\qquad\times\begin{pmatrix}F&1&F^{\prime}\\ m_{F}&q&-m_{F}^{\prime}\end{pmatrix}^{2}.\end{split}$ (S13)
The right-hand side of this equation is proportional to the squared right-hand side of (S12), with the notable feature that, for a given initial state, the sum over all $q$ and possible jumps is unity. Thus the path’s $\gamma_{k}$ (which is shared for all $q$) is the total radiative lifetime for each initial state.
We treat decay from the Rydberg states as primarily incoherent. The vast majority of possible radiative decay paths pass through states outside of those included in our model. The long lifetime of the Rydberg states also means that the rates of non-radiative incoherent decay processes are of similar magnitude to the radiative decay. These include transit-time broadening [A. Elliott et al., in preparation] and collisional broadening [65]. We note the rudimentary model for transit-time broadening [66] is only valid for large beams where the decay rate is larger than the transit rate.
The incoherent decay paths are included in our model by means of a “dummy state” with no angular momentum character. Excited states decay to this dummy state via a collapse operator with a single non-zero element. In the case of the $i$-level states, the rate is given by the sum of the transit-time broadening and collisional broadening rates. For the two Rydberg levels, the radiative decay rate is added to this. The “dummy state” itself decays rapidly into the $|g\rangle$ manifold (with a near-infinite $\gamma$).

### v5

@@ IV Numerical simulations
where $I_{0}$ is the intensity of the probe light before traversing the distance $\ell$ through the vapour. The extinction coefficient $\alpha$ is proportional to a weighted sum over the off-diagonal elements of the density matrix that involves the combination of $g$ and $i$ -states [See Eq. (S15) of Supplementary Note 2]. To simulate the variation of $\alpha(\theta,\Delta_{\text{c}})$ over a spectrogram for a given RF dressing power, we therefore calculate the steady-state atomic density matrix under the combined influence of the optical and RF fields (see appendix C). The density matrix for each case includes all individual $|Fm_{F}\rangle$ hyperfine states of the ladder levels (cf. Fig. 2 and Fig. 3). For the type-1 ladder of Fig. 2 involving a $\mathrm{D}_{5/2}\!\!\leftrightarrow\!\!\mathrm{P}_{3/2}$ Rydberg transition, this amounts to 52 atomic states, whereas the type-II ladder of Fig. 3 capped by the $35\mathrm{D}_{3/2}\!\!\leftrightarrow\!\!36\mathrm{P}_{1/2}$ transition requires 36 atomic states. As in Ref. 34 an auxiliary “dummy state” is included to incorporate effective incoherent decay from transit-time broadening, and radiative decay from the Rydberg states to states not explicitly included in the model. In our simulations, the RF power is the only fitted quantity and other input parameters to the modelling are either directly measured as in the case of optical beam waists and powers or they are tabulated values, as in case of 87Rb atomic properties. Figure 8 showcases our ability to reliably simulate spectrograms for both type-I and type-II ladders.
@@ V Discussion
Figure 9: Electrometry using a type-I system with co-polarized fields. (a) Experimentally observed AT splitting of the $34D_{5/2}$ level centred on $\Delta_{\text{c}}=0$. The feature at $\Delta_{\text{c}}\sim-306$ MHz is the $34D_{3/2}$ level [cf. Fig. 6(c)]. Interaction with this state influences the $\Delta_{\text{c}}<0$ AT peak. (b) Development in the $\Delta_{\text{c}}>0$ AT peak with RF power increasing (top to bottom) from 10 to 80 (arb. units) in uniform steps. (c) Simulated spectrogram of the probe field transmission with artificially reduced broadening. (d) Simulated probe field transmission for co-polarized fields ($\theta=180^{\circ}$) with artificially reduced broadening.
Figure 10: Level diagrams for a $J=\frac{1}{2}\frac{3}{2}\frac{5}{2}\frac{3}{2}$ system (a) and a $J=\frac{1}{2}\frac{3}{2}\frac{3}{2}\frac{1}{2}$ system (b) driven by $\pi$-transitions. Both realise two parallel ideal four level ladders (green background). (c) $m_{F}=1$ ladder of Fig. 2(b) in isolation. The system cannot be broken down into parts as shown in (d). (e) Level splittings obtained from diagonalizing the matrix describing the RF-coupling (orange arrow) amongst the $F$-states of the Rydberg levels $r_{1}$ and $r_{2}$ when $m_{F}=1$. The different $F$-parts of $r_{1}$ and $r_{2}$ have been colour-coded to visualise their contributions to each hybridized dressed state. The diagonalization for the $m_{F}$=1 ladder yields five eigenenergies with separations and multiplicities as annotated.
For an ideal four-level ladder system, the AT-splitting that emerges in the EIT spectrum is directly proportional to the amplitude of the RF-field dressing the topmost two Rydberg states and the constant of proportionality is given by a combination of fundamental constants and the transition dipole moment between the Rydberg states. Because of the quasi-one-electron character of Rydberg atoms, the latter can be calculated to a very high accuracy. Early work, including Ref. 12, however, has introduced the misconception that for a type-I ladder with linear co-polarized fields, one is dealing with a single transition dipole moment between $r_{1}$ and $r_{2}$, and this has been reiterated pervasively in the literature [36, 37, 38, 39, 40, 41, 14, 42, 43, 44, 45, 46, 47]. In fact, two transition dipole moments are in play differing by a factor of $\sqrt{2/3}\approx 0.8$ [cf. Fig. 4(a)] and, as noted in our above analysis, one will not be measuring a pure AT-doublet but a four-peaked spectrum. We illustrate this effect in Fig. 9(a) which presents the development in AT splitting when moving our RF emitter up close ($\sim 2$ cm) to the vapor cell (see appendix B). Focusing our attention to the blue-detuned branch ($\Delta_{c}>0$), we se how the split-out peak becomes lopsided [see Fig. 9(b)] for high RF powers. If the goal is to establish an accurate electrometry standard based on a type-I ladder system, the four-peak nature should be accounted for because the splitting of each doublet-pair tunes at different rates: the inferred field amplitude will be skewed if an unresolved four-peak spectrum is interpreted as a pure doublet. Figure 9(c) shows a simulated spectrogram for a type-I system with Doppler broadening “turned off” from which Fig. 9(d) picks out the parallel-fields-spectrum that underpins the skewing of the experimental high power trace in Fig. 9(b).
The spectrogram in Fig. 9(c) also elucidates that for linearly-polarized fields, the eigenenergy corresponding to a split-out peak does not change with $\theta$: the spectral features are straight vertical lines. While the position of a peak remains fixed, its prominence in the spectrum will, however, generally change. For two overlapping broadened peaks, this may give the appearance of single feature that moves continuously as $\theta$ changes. This effect can be observed as wobbling of the side-lobes in the type-1 column of Fig. 8.
@@
To be clear, it is incorrect to talk about “the” dipole moment between $D_{5/2}$ and $P_{3/2}$ levels for a linearly-polarized coupling field [48, 42, 43, 47, 49, 50, 51] because it depends on whether $m_{J}=|1/2|$ or $m_{J}=|3/2|$. It is also incorrect to resort to using the $m_{J}=|1/2|$ value [36, 45, 38, 39, 52], based on a rationale of all fields being $\pi$-polarized. This would be applicable in the absence of hyperfine structure of the ground state [see Fig. 10(a) and (b)], in which case the system is described by two parallel ideal four-state systems. For an atom in a specific, resolved hyperfine state we are, however, not dealing with parallel four-state systems, even when driving with co-polarized fields. This is evident from the diagrams in Figs. 2(b) and 3: moving up from any $m_{F}$-state of the ground level one counts more than three transition-arrows. Moreover, a vertical cascade corresponding to a given $m_{F}$ [Fig. 10(c) shows the example of $m_{F}=1$ from Fig. 2(b)] cannot be broken down into individual “unique contributions” [cf. Fig. 10(d)] as was recently stated in Ref. 53. The claim of Ref. 53 that the EIT scheme acts to probe the states of $r_{1}$ individually is invalidated by the fact that these states are degenerate and therefore indirectly coupled with each other by the RF-field via the states of $r_{2}$ [54]. Instead, when treating the the example of the $m_{F}=1$ cascade in the hyperfine basis, the splittings of the system probed via EIT is found by diagonalizing the entire $7\times 7$ matrix that describes the resonant RF interactions between the hyperfine states of $r_{1}$ and $r_{2}$ . The result of this procedure is illustrated in Fig. 10(e), replicating the $m_{F}=1$ dressed energies encountered in our procedure Fig. 4c where the eigenenergies are first found in $J$-basis basis after which the system is recast in the $F$-basis.
@@
The insight that the co-polarized cases of Figs. 2(b) and 3 are not represented by a collection of four-level systems means that there is a priori no reason why they should give rise to simple AT doublets. Indeed, as we have shown, the type-4 system features a prominent central peak in its spectrum. Interestingly, the two sidelobes of the type-4 triplet tunes as an ideal AT doublet with RF power with a tuning rate described by a single transition matrix element. The perfect absence of a central peak of the $\theta=0^{\circ}$ type-1 system in Fig. 2(b) can be interpreted as a destructive interference between the multiple parallel pathways represented by Eq. (3). Regardless, the spectrum does not represent a simple AT doublet.
@@ VI Conclusion and outlook
In the current study, we have focused on the case of linearly-polarized optical and RF fields and we have shown that either of the two complementary spectroscopic signatures for type-1 and -4 ladders could serve to determine the angle of polarization for the RF field. The simultaneous high fidelity of our density matrix simulations for these complementary test cases instils confidence in our simulation framework’s ability to synthesize the probe light transmission for more complex combinations of polarization states. In the future we plan to pursue the determination of arbitrary polarization states of the incoming RF field through machine learning approaches where the polarization states of probe and coupling light are varied to acquire a sequence of spectrograms. For this purpose the data diversity [56] provided by the profoundly different response from type-1 and type-4 may provide a significant advantage. This would open up a new paradigm for quantum-metrological [57] polarimeters that at its heart is enabled by the quantisation of atomic angular momentum.
@@ Acknowledgements.
M. Cloutman performed experiments with support from M. Chilcott, A. B. D., and S. O.; M. Cloutman analysed the data with support from M. Chilcott; M. Chilcott and A. E. carried out density matrix calculations; N. K., M. Cloutman, and M. Chilcott wrote the manuscript with input from all authors; N. K. conceived and supervised the project.
@@ Appendix B RF dressing field
The $\sim 55$ GHz RF field dressing our Rydberg atoms was produced by a photomixing antenna (Toptica Tera-Beam 1550). Two $\sim 1550\text{\,}\mathrm{n}\mathrm{m}$ fiber-coupled DFB lasers are combined to produce a beatnote with frequency $f_{\text{RF}}$ on an InGaAs photoconductive chip mounted on a bowtie-shaped antenna. The photomixing process produces linearly-polarized radiation at $f_{\text{RF}}$, which is collimated to a beam diverging at $12\text{\,}\mathrm{\SIUnitSymbolDegree}$ by a silicon lens. The frequencies of the DFB lasers are temperature tuned, allowing us to target the the $34\mathrm{D}_{5/2}\!\!\leftrightarrow\!\!35\mathrm{P}_{3/2}$ and $35\mathrm{D}_{3/2}\!\!\leftrightarrow\!\!36\mathrm{P}_{1/2}$ transitions at $56.714\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$ and $53.840\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$ respectively. The emitter is positioned $15\text{\,}\mathrm{c}\mathrm{m}$ above the vapour cell which is mounted on a wooden post to minimise metallic interference. Running the experiment using a $S_{1/2}\leftrightarrow P_{1/2}$ transition to verify that this gives rise to a spectrogram with the two straight horizontal lines [21] we ascertain that RF-radiation reflected back on the cell is insignificant: reflected RF radiation field will change the polarization state at the atoms and bend the lines in the spectrogram [22].
For the measurements presented in Figure 9(b) and (c), the emitter was positioned very close ($2\text{\,}\mathrm{cm}$) to the cell. A $0.75\text{\,}\mathrm{mm}$-pitch wire grid polarizer was mounted in between the emitter and the cell to suppress effects a complex RF field polarization [22] and ensure a linear polarization at the position of the atoms.
@@ Appendix C Density matrix calculations
Here the polarizations and Rabi frequencies of the probe, coupling and RF fields are encoded in the coupling operators $\Omega_{k}$ between manifolds, represented in (7) using an ordered basis of the $g$, $i$, $r_{1}$ and $r_{2}$ levels. The detuning blocks $\delta_{k}$ are scalar matrices because the individual states acted on by each block are not resolved. The problem is described in the $|Fm_{F}\rangle$ basis, to deal with the fact that the probe laser field resolves transitions between hyperfine states. The second, dissipative term of Eq. (6) is composed of collapse operators $L_{i}$ with corresponding decay rates $\gamma_{i}$. A representation of $L_{i}$ for radiative decay in the $|Fm_{F}\rangle$ basis is given in Supplementary Note 2. This term facilitates the inclusion of two classes of state decay: spontaneous emission, which has an operator for each possible polarization of the radiated photon, and incoherent decay, which is modelled with an axillary “dummy state” [34]. Incoherent decay includes transit-time broadening of all states, but it also describes some radiative decay of the Rydberg states, which usually involves the emission of multiple photons on a decay path outside the state space of the simulation. The “dummy state” then decays rapidly, with equal probability, to each sublevel of the ground-state manifold.
@@
While the counter-propagating arrangement of coupling and probe beams compensates for the Doppler effect, we include the residual Doppler broadening of the room-temperature vapor by performing simulations with detunings that correspond to a range of atomic velocities along the probe/coupling beams, and averaging them according to the Maxwell-Boltzmann distribution. Our lock-in detection scheme is accounted for by removing a constant offset in the simulation: the resonant extinction of the probe beam without the coupling field.
@@
To arrive at Eq. (3) we must evaluate $\left|\sum_{F=1}^{4}C^{Fm_{F}}_{Jm_{J}Im_{I}}\left<F^{\prime}m_{F}^{\prime}\right|r_{0}\left|Fm_{F}\right\rangle\right|^{2}$.
@@
In this supplement, we provide additional details on computing the density matrix, and connecting it to experimental observations. In particular, we include the matrix form of the operators necessary to formulate the master equation Eq. (6), expanded over the hyperfine $|Fm_{F}\rangle$ basis. The Python computer code used to produce the simulated results of Fig. 8 are available as part of this submission.
@@
The coupling operator $\Omega_{k}$ for each field ($k=p,c,\text{RF}$) is given by,
@@
where $\vec{E}_{k}$ is the electric field, and $\vec{\mu}_{k}$ is a vector operator describing the dipole moments between states driven resonantly.
To populate the coupling matrix Eq. (7), we make use of a spherical basis expansion [66]. The spherical basis components $\hat{E}^{(1)}$ of the unit vector along the electric field direction (i.e., $\vec{E}=|\vec{E}|\hat{E}$), is given in terms of the Cartesian components by
$\hat{E}^{(1)}$ $=-\frac{1}{\sqrt{2}}(\hat{E}_{x}+\text{i}\hat{E}_{y}),$ (S7)
$\hat{E}^{(0)}$ $=\hat{E}_{z},$ (S8)
$\hat{E}^{(-1)}$ $=\frac{1}{\sqrt{2}}(\hat{E}_{x}-\text{i}\hat{E}_{y}).$ (S9)
Physically, q=1,0,+1q=-1,0,+1 represents the change in angular momentum associated with $\sigma^{-},\pi,$ and $\sigma^{+}$ excitations, respectively. Similarly, we expand $\vec{\mu}_{k}$ over the spherical basis $\hat{\epsilon}_{q}$, to introduce the angular component matrices $u_{k}^{(q)}$,
$\vec{\mu}_{k}\propto\sum_{q=-1,0,1}\hat{\epsilon}_{q}u_{k}^{(q)}.$ (S10)
Using the spherical basis expansions, we factor the coupling operators into a scalar radial component $\Omega^{(r)}_{k}$ (the radial Rabi frequency) and an angular component,
$\Omega_{k}=\Omega_{k}^{(r)}\sum_{q=-1,0,1}\hat{E}^{(q)}_{k}u_{k}^{(q)},$ (S11)
where the Radial rabi frequency between levels of principle quantum numbers $n$ and $n^{\prime}$, and orbital angular momenta $L$ and $L^{\prime}$ is
$\Omega^{(r)}_{k}$ $=|\vec{E}|e\int_{0}^{\infty}r^{2}\text{d}r\;\psi_{n^{\prime}L^{\prime}}^{*}(r)r\psi_{nL}(r),$ (S12)
in terms of the radial wavefunctions $\psi(r)$. We calculate this integral using the ARC [58].
In the hyperfine basis, the elements of each matrix $u_{k}^{(q)}$ in Eq. (S11) are,
@@
$\begin{split}&\left\langle SLJFm_{F}\right|L_{k}^{(q)}|S^{\prime}L^{\prime}J^{\prime}F^{\prime}m_{F}^{\prime}\rangle^{2}=\\ &\qquad(2L+1)(2J+1)(2J^{\prime}+1)(2F+1)(2F^{\prime}+1)\\ &\qquad\times\begin{Bmatrix}L^{\prime}&J^{\prime}&S\\ J&L&1\end{Bmatrix}^{2}\times\begin{Bmatrix}J^{\prime}&F^{\prime}&I\\ F&J&1\end{Bmatrix}^{2}\\ &\qquad\times\begin{pmatrix}F^{\prime}&1&F\\ -m_{F}^{\prime}&q&m_{F}\end{pmatrix}^{2}.\end{split}$ (S14)
The right-hand side of this equation is proportional to the squared right-hand side of Eq. (S13), with the notable feature that, for a given initial state, the sum over all $q$ and possible jumps is unity, so the decay path’s $\gamma_{k}$ is the total radiative rate.
In our system, the decay from the Rydberg states is dominated by the non-radiative transit-time broadening [67] with a rate $\gamma_{\text{TTB}}\approx 2\pi\times$200\text{\,}\mathrm{kHz}$\$. Additionally, the vast majority of possible radiative Rydberg decay paths pass through states not included in our model. These (effectively) incoherent decay paths are modelled by means of a “dummy state” with no angular momentum character. Excited states decay to this dummy state via a collapse operator with a single non-zero element. In the case of the $i$-level states, the rate is given by the the transit-time broadening. For the two Rydberg levels, the total radiative decay rate to states other than $\left|i\right$> is added to this. The “dummy state” itself decays rapidly into the $|g\rangle$ manifold (with a near-infinite $\gamma$). We note the rudimentary model for transit-time broadening used is valid for large beams where the radiative decay rate is larger than the transit rate, and this system’s dominant radiative decay process is the $\gamma_{i}\approx 2\pi\times$6\text{\,}\mathrm{MHz}$\$.
Probe extinction
As discussed in the main text, the extinction of the probe beam of polarization $q$ through an atomic vapour follows the Beer-Lambert law, Eq. (5). The extinction coefficient for a $q$-polarized field, $\alpha^{(q)}$, through a vapour of density $n$ at a transition of frequency $\omega_{p}$ is composed of off-diagonal elements of the atomic density matrix $\rho$,
$\alpha^{(q)}=\frac{2n\omega_{p}}{c\epsilon_{0}\hbar}\sum_{g_{l}\in g,i_{m}\in i}\left(\mu^{(q)}_{i_{m}g_{l}}\right)^{2}\text{Im}\left(\frac{\rho_{g_{l}i_{m}}}{\Omega_{g_{l}i_{m}}}\right),$ (S15)
with the sum running over the $q$-coupled states of the $g$- and $i$-level manifolds. The density matrix is the steady state of the master equation of Eq. (6) with the atomic Hamiltonian Eq. (7), and the operators described above.